OpenStudy (anonymous):

Solve: 0 = 6x2 − 10x − 4. x = −1/3, x = −2 x = −1/3, x = 2 x = 1/3, x = −2 x = 1/3, x = 2

OpenStudy (anonymous):

are you just looking for the answer or are you trying to learn the material?

OpenStudy (anonymous):

learn

OpenStudy (anonymous):

okay, do you know the quadratic equation?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

take a look here: http://www.mathsisfun.com/algebra/quadratic-equation.html

OpenStudy (anonymous):

ok so the 2 squared makes it quadratic

OpenStudy (anonymous):

the x^2 makes it quadratic, right. The general form is ax^2 + bx + c = 0. In your problem, what does that make the values of a, b, and c?

OpenStudy (anonymous):

ax^2 + bx + c = 0 6^2 - 10x - 4 = 0

OpenStudy (anonymous):

well you can't take the X out just yet, but just by inspecting the equation you should be able to see the values of A, B, and C. For example, C is -4 since it's the term without an X. can you tell me what A and B are?

OpenStudy (anonymous):

start with 6x^2 - 10x - 4 = 0

OpenStudy (anonymous):

A= 6x^2 B= -10x

OpenStudy (anonymous):

a, b, and c are the terms that appear in front of the x^2 and x terms. For example, if the equation were 9x^2 + 4x - 8, then a=9 b=4 c=-8 So, given that example, what numbers are A, B and C for your equation?

OpenStudy (anonymous):

a=6 b=-12 c=-4

OpenStudy (anonymous):

close -- you got A and C right, but not B, maybe just a typo?

OpenStudy (anonymous):

ya srry -10

OpenStudy (anonymous):

right! okay, now you have A, B and C so you can just plug them into the quadratic formula to get the two values of X that make the equation true. it's this: http://www.mathsisfun.com/algebra/images/quadratic-formula.gif

OpenStudy (anonymous):

there may be two solutions, which you can find by using the "+" in the formula one time, and the "-" in the other time (where it says +/-)

OpenStudy (anonymous):

i dont get this part

OpenStudy (anonymous):

okay well you start with the quadratic formula, and every place you see A, B or C you substitute in the numbers for A B and C for your problem

OpenStudy (anonymous):

_______________ x = -10 + / 10^2 - 46-4

OpenStudy (anonymous):

_______________ x = -10 - / 10^2 - 46-4

OpenStudy (anonymous):

well the first term in the top left is -b, and since for you b=-10, the negatives cancel so it's just 10

OpenStudy (anonymous):

and (4)(a)(c) is 4 * 6 * -4

OpenStudy (anonymous):

_______________ x = 10 + / 10^2 - 46-4

OpenStudy (anonymous):

remember, "4ac" means 4 * a * c

OpenStudy (anonymous):

_______________ x = 10 + / 10^2 - 4*6*4

OpenStudy (anonymous):

and remember the division, it's all divided by 2a

OpenStudy (anonymous):

then compute that and you'll get one of the two answers

OpenStudy (anonymous):

2*6? 12

OpenStudy (anonymous):

right, so put it all together and what answer do you get?

OpenStudy (anonymous):

10+/10^2 -96

OpenStudy (anonymous):

well remember 4 a c has a negative term in the c

OpenStudy (anonymous):

so when you multiply it all out you get a negative number, which cancels with the negative in the equation

OpenStudy (anonymous):

so just 96

OpenStudy (anonymous):

right, so what's b^2 - 4 a c?

OpenStudy (anonymous):

10^2 + 96

OpenStudy (anonymous):

right, what does that add up to?

OpenStudy (anonymous):

106^2

OpenStudy (anonymous):

no...what's 10^2?

OpenStudy (anonymous):

oh 100

OpenStudy (anonymous):

so 196

OpenStudy (anonymous):

right so what's 10^2 + 96?

OpenStudy (anonymous):

and that's under a square root, so what's the square root of 196

OpenStudy (anonymous):

14

OpenStudy (anonymous):

yep so can you solve the whole quadratic formula now?

OpenStudy (anonymous):

10 + / 14 -------- ??????????????? 12

OpenStudy (anonymous):

(10 + 14)/12 and (10 - 14)/12 are the two answers

OpenStudy (anonymous):

note that the quadratic formula has a "+ or -" before the square root, meaning you do it once with + and a second time with - to get the two answers

OpenStudy (anonymous):

ok so 24/12=1/2 and -4 / 12 = -3

OpenStudy (anonymous):

close -- 24/12 = 2, not 1/2

OpenStudy (anonymous):

the second answer you made the same mistake

OpenStudy (anonymous):

oh sorry lol

OpenStudy (anonymous):

i got 0.3333...

OpenStudy (anonymous):

yeah, -1/3 is the right answer

OpenStudy (anonymous):

ok so -1/3 and 2

OpenStudy (anonymous):

check your multiple choices

OpenStudy (anonymous):

yep

Parth (parthkohli):

Why the quadratic formula?

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

because it works in every case

OpenStudy (anonymous):

yep, it's B

Parth (parthkohli):

\[6x^2 - 10 x - 4 = 6x^2 - 12x + 2x - 4 = 6x(x - 2) + 2(x - 2) = (6x + 2)(x - 2)\cdots\]

Parth (parthkohli):

Now the zero-product property.

OpenStudy (anonymous):

why do that instead of the quadratic formula?

Parth (parthkohli):

Because it is simpler

OpenStudy (anonymous):

i disagree